OpenStudy (anonymous):

Help me plz y'''+6y''+y'-34y=0 if y1=e^-4x

OpenStudy (anonymous):

What do you mean by y1? @RaphaTkt

OpenStudy (anonymous):

I want to know how to find the other two roots by the method of reducing job order

OpenStudy (amistre64):

doesnt this have to do with a wronskian? i tend to get the methods mixed up

OpenStudy (anonymous):

I say just find the two missing roots of ED

OpenStudy (anonymous):

\[y'''+6y''+y'-34y=0\] The characteristic equation is \[r^3+6r^2+r-34=0\] The given solution \(y_1=e^{-4x}\) tells you that \(r=-4\) is a root of the characteristic equation. This means that you can write the characteristic polynomial as \[(r+4)(r^2+\cdots)=0\] To find the \(r^2+\cdots\) part I think you'll have to do some long division: \[\frac{r^3+6r^2+r-34}{r+4}=r^2+2r-7-\frac{6}{r+4}\] However, checking with WA, \(r=-4\) is not an root of the polynomial: http://www.wolframalpha.com/input/?i=r%5E3%2B6r%5E2%2Br-34%3D0 In fact, the solution doesn't contain a term of only \(e^{-4x}\): http://www.wolframalpha.com/input/?i=y%27%27%27%2B6y%27%27%2By%27-34y%3D0